OpenStudy (tiffany_rhodes):

Prove that a normal operator with real eigenvalues is self-adjoint. I know a matrix (operator) is normal if: N*N = NN*. Also, I know a matrix is self-adjoint (hermitian) if N = N*. Is the proof as easy as: (N*N)* = N*N** = N*N=NN* (properties of adjoints and definition of normal). (NN*)* = N**N* = NN* = N*N ? It seems a little too easy. It doesn't seem like I'm proving N = N*

OpenStudy (stamp):

Wew! This is a spicy one.

OpenStudy (stamp):

I have never studied this subject in-depth but after some Google-Fu this looks useful: https://www.math.ucdavis.edu/~anne/WQ2007/mat67-Ll-Spectral_Theorem.pdf Good luck.

OpenStudy (518nad):

you just stated what a normal operator is

OpenStudy (loser66):

The normal operator T is self adjoint with real eigenvalues

OpenStudy (loser66):

Because \[M(T) = \left[\begin{matrix}\lambda_1 &0&0&\cdots\\0&\lambda_2&0&\cdots\\\cdots\\\cdots &0&\cdots&\lambda_n\end{matrix}\right]\]

OpenStudy (loser66):

T is self adjoint iff T = T* but \[(T)^* = \left[\begin{matrix}\bar\lambda_1 &0&0&\cdots\\0&\bar\lambda_2&0&\cdots\\\cdots\\\cdots &0&\cdots&\bar\lambda_n\end{matrix}\right]\]

OpenStudy (loser66):

In \(\mathbb R\), \(\lambda =\bar\lambda\)

OpenStudy (loser66):

So, T is self-adjoint. Dat sit

OpenStudy (loser66):

Sorry, \(M(T) = M(T^*)\), lack of M in the front. :)

OpenStudy (518nad):

okay i think i havee a proof for u